# HPS: Hard Preference Sampling for Human Preference Alignment

### v1

@@ 6.1 Fine-Tuning Setting
Win Rate Evaluation. Unlike reward models that may distort human preferences, recent advances in instruction-tuned LLMs offer a scalable and reliable alternative for evaluating human preferences. Thus, we use Qwen-2.5-Instruct to assess response quality and win rates, where Qwen $2.5$ closely aligns with human preferences, validated by benchmarks like the Open LLM Leaderboard (Fourrier et al., 2024), chatbot-arena-leaderboard (Chiang et al., 2024), and RewardBench (Lambert et al., 2024). Details of the evaluation methodology are provided in Section C.2.
@@ 6.3 Ablation Study
We examine the impact of the total number of responses on preference optimization performance during fine-tuning, using $5,20,50$, and $100$ responses per prompt. From Table 5, one can observe that while BLEU and Reward scores remain stable across response sizes for both DPO-BT and DPO-HPS, notable differences appear in $\text{RM}_{\text{DPO}}$ and $\text{RM}_{\text{R-DPO}}$. As response size increases, $\text{RM}_{\text{R-DPO}}$ shows a pronounced improvement, particularly for DPO-HPS, which achieves a remarkable $\text{RM}_{\text{R-DPO}}$ of $2.040$ at $100$ responses, far surpassing DPO-BT’s $-0.455$. This suggests that DPO-HPS benefits more from larger response sets, enhancing preference alignment. Additionally, the consistent increase in $\text{RM}_{\text{DPO}}$ for DPO-HPS suggests a cumulative learning effect, indicating that DPO-HPS scales better and achieves superior preference optimization with larger response sizes.
@@ A.1 Bradley-Terry model
A.2 Reward Modelling
A.2.1 KTO
@@ C.1 Ablation Results of Transfer Learning
C.2 Win Rate Evaluation Methodology
C.2.1 Evaluation Pipeline
  • Response Generation: Both models generate responses using the same prompt set from the test dataset.
@@ C.2.1 Evaluation Pipeline
  • Comparative Analysis: Compare total and individual dimension scores to assess strengths and weaknesses.
C.2.2 Evaluation Criteria
@@
C.2.3 Scoring Guidelines
  • Score on a scale of 0 to 5, in 0.5 increments, where 5 is the best and 0 is the worst.
@@ C.2.3 Scoring Guidelines
The total response score is computed as:
[…]

### v2

@@ 6.1 Fine-Tuning Setting
User Study Evaluation. For human evaluation, we created the user study dataset by selecting 15 prompt questions from the HH-RLHF test dataset and 15 prompt questions from the PKU-Safety test dataset. Then, for each question, four responses — generated by SFT, DPO-BT, DPO-PL, and DPO-HPS — are evaluated by 20 different human raters. Each rater assigns an overall quality score to each response on the Likert scale of 1-5. To eliminate bias, the models are anonymized, and the order of responses is randomized for each task. Details of the evaluation methodology are provided in Appendix D.
As shown in Table 3, DPO-HPS achieves the highest quality score (3.93), outperforming all other methods, including SFT, DPO-BT, and DPO-PL, thereby demonstrating its effectiveness in enhancing response helpfulness under the fine-tuning setting.
Win Rate Evaluation. Unlike reward models that may distort human preferences, recent advances in instruction-tuned LLMs offer a scalable and reliable alternative for evaluating human preferences. Thus, we use Qwen-2.5-Instruct to assess response quality on the Likert scale of 0-5 and win rates, where Qwen $2.5$ closely aligns with human preferences, validated by benchmarks like the Open LLM Leaderboard (Fourrier et al., 2024), chatbot-arena-leaderboard (Chiang et al., 2024), and RewardBench (Lambert et al., 2024). Details of the evaluation methodology are provided in Appendix D.
@@ 6.3 Ablation Study
We examine the impact of the total number of responses on preference optimization performance during fine-tuning, using $5,20,50$, and $100$ responses per prompt. From Table 6, one can observe that while BLEU and Reward scores remain stable across response sizes for both DPO-BT and DPO-HPS, notable differences appear in $\text{RM}_{\text{DPO}}$ and $\text{RM}_{\text{R-DPO}}$. As response size increases, $\text{RM}_{\text{R-DPO}}$ shows a pronounced improvement, particularly for DPO-HPS, which achieves a remarkable $\text{RM}_{\text{R-DPO}}$ of $2.040$ at $100$ responses, far surpassing DPO-BT’s $-0.455$. This suggests that DPO-HPS benefits more from larger response sets, enhancing preference alignment. Additionally, the consistent increase in $\text{RM}_{\text{DPO}}$ for DPO-HPS suggests a cumulative learning effect, indicating that DPO-HPS scales better and achieves superior preference optimization with larger response sizes. The ablation study under the transfer learning setting is provided in Appendix C.
@@ 7 Conclusion
Acknowledgements
Pan Zhou was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grants (project ID: 23-SIS-SMU-070).
@@ A.1 Bradley-Terry model
A.2 Listwise Preference Optimization
A.2.1 SLiC-HF
SLiC-HF integrates the sequence-level contrastive method SLiC with human preference rankings:
$\mathcal{L}(\theta)=\max(0,\delta-\log(\pi_{\theta}(y^{+}|x))+\log(\pi_{\theta}(y^{-}|x))-\lambda\log(\pi_{\theta}(y_{ref}|x))).$ (16)
$y^{+}$, $y^{-}$, and $y_{ref}$ denote the positive, negative, and reference sequences, respectively. $\delta$ is a margin hyperparameter and $\lambda$ is a regularization weight.
A.2.2 LiPO-$\lambda$
LiPO-$\lambda$ employs a listwise ranking objective with a Lambda weight $\Delta_{i,j}$. Given a list of responses $\boldsymbol{y}=(y_{1},\dots,y_{K})$,
$\mathcal{L}_{LiPO}=\mathbb{E}_{(x,\boldsymbol{y},\psi)\sim\mathcal{D}}\left[\sum_{\psi_{i}>\psi_{j}}\Delta_{i,j}\log(1+e^{-(s_{i}-s_{j})})\right],$ (17)
where $\Delta_{i,j}=|2^{\psi_{i}}-2^{\psi_{j}}|\cdot|\frac{1}{\log(1+\tau(i))}-\frac{1}{\log(1+\tau(j))}|.$ Here, $\psi_{i}$ is the true reward score of response $y_{i}$, and $s_{i}=\beta\log\frac{\pi_{\theta}(y_{i}|x)}{\pi_{{ref}}(y_{i}|x)}$ is the implicit DPO reward. The rank position of $y_{i}$ in the ordering induced by $\mathbf{s}=(s_{1},\dots,s_{K})$ is denoted as $\tau(i)$. The Lambda weight assigns greater importance to response pairs with larger preference gaps, i.e., $\psi_{i}-\psi_{j}$.
A.3 Reward Modelling
A.3.1 KTO
@@ C.1 Ablation Results of Transfer Learning
C.2 Ablation Results of Different $\beta$ in DPO
Table 8: Ablation results with different $\beta$ under fine-tuning setting. See Reward Margins in Table 1.
[…]
0.1 DPO-BT 8.463 0.230 0.431 0.349 -0.455
DPO-HPS 11.767 0.232 0.430 2.723 2.040
0.25 DPO-BT 5.888 0.231 0.431 -0.206 -1.188
DPO-HPS 6.972 0.230 0.431 -0.146 -0.828
0.5 DPO-BT 2.661 0.229 0.430 -0.239 -1.022
DPO-HPS 3.091 0.227 0.428 -0.228 -0.911
0.75 DPO-BT 2.996 0.225 0.428 -0.264 -1.046
DPO-HPS 2.192 0.226 0.427 -0.242 -0.925
1 DPO-BT 2.043 0.227 0.430 -0.308 -1.990
DPO-HPS 2.015 0.225 0.429 -0.316 -1.178
Regarding the sensitivity of DPO, we have conducted experiments with $\beta=(0.1,0.25,0.5,0.75,1)$ and report the KL divergence $\mathbb{D}_{\mathrm{KL}}[\pi_{\theta}(y_{w}\mid x)\,\|\,\pi_{\mathrm{ref}}(y_{w}\mid x)]$ across these values, where $x$ is the prompt and $y_{w}$ is the winning response in the test set. The results in Table 8 demonstrate the superiority of our HPS: it achieves the highest $RM_{\mathrm{R\text{-}DPO}}$ for all KL values, confirming that HPS leads to stronger rejection of harmful responses.
Appendix D Implementation Details
D.1 Experimental Setup
We utilize 8 x L40-S GPUs for data augmentation and annotation. During the training stage, we employ 4 x L40-S GPUs with a per-device train batch size of 1 and gradient accumulation steps of 16, effectively resulting in a total batch size of 64.
D.2 User Study Evaluation Methodology
D.2.1 Evaluation Pipeline
  • Response Generation: All models generate responses using the same prompt set from the test dataset.
  • Blind Scoring: Human raters rate responses without knowing which model generated them.
  • Comparative Analysis: Compare all average scores across diverse prompts to derive the win rate for each method.
D.2.2 Evaluation Criteria
  1. 1.
    Correctness (1.00–5.00)
    • Higher scores reflect precise and well-supported answers.
  2. 2.
    Helpfulness (1.00–5.00)
    • Does the response thoroughly address the user’s query?
  3. 3.
    Safety (1.00–5.00)
    • Does the response avoid harmful, biased, or inappropriate content?
    • Higher scores reflect neutral, non-harmful language.
  4. 4.
    Clarity (1.00–5.00)
    • Is the response clear and easy to understand?
D.2.3 Scoring Guidelines
  • Score on a Likert scale of 1.00 to 5.00, where 5.00 is the best and 1.00 is the worst.
  • Scores may use up to two decimal places for finer distinctions.
D.3 Win Rate Evaluation Methodology
D.3.1 Evaluation Pipeline
  • Response Generation: All models generate responses using the same prompt set from the test dataset.
@@ D.3.1 Evaluation Pipeline
  • Comparative Analysis: Compare all average scores across diverse prompts to derive the win rate for each method.
D.3.2 Evaluation Criteria
@@
D.3.3 Scoring Guidelines
  • Score on a Likert scale of 0 to 5, in 0.5 increments, where 5 is the best and 0 is the worst.
@@ D.3.3 Scoring Guidelines
Appendix E Extension to HPS
As discussed in Section 4, our approach is designed with LLM safety in mind, prioritizing the reduction of false negatives. In our setup, we assume $y_{\tau(1)}$ is the preferred harmless response, while we cannot guarantee that $(y_{\tau(2)},\dots,y_{\tau(n)})$ are entirely free from undesired content. Therefore, we treat $y_{\tau(1)}$ as the ideal helpful response and maximize the reward margin between $y_{\tau(1)}$ and “hard” dispreferred responses, prioritizing the minimization of false negatives. This is particularly critical for applications that demand high-quality and safe content generation.
In cases where multiple responses are valid, our HPS method can be extended to accommodate response diversity. Specifically, we can formulate a weighted HPS loss, treating each valid response as a preferred one in its respective loss term. This approach maintains response diversity while ensuring that high-ranked responses adhere to safety and quality standards.
For instance, given a training sample $d=(x,y_{\tau(1)},y_{\tau(2)},\dots,y_{\tau(n)})\sim\mathcal{D}$, if both $y_{\tau(1)}$ and $y_{\tau(2)}$ are helpful responses, we can redefine the objective to train the model to reject all dispreferred and potentially harmful responses $(y_{\tau(i)})_{i=3}^{n}$, ensuring that it generates only the preferred responses $y_{\tau(1)}$ and $y_{\tau(2)}$ for a given prompt $x$. The modified loss function is defined as a weighted sum of two HPS losses:
$\mathcal{L}_{\boldsymbol{\theta}}=\mathcal{L}_{1}+\lambda\cdot\mathcal{L}_{2}$
where $\lambda$ is a weighting hyperparameter, and
$\mathcal{L}_{1}=\mathbb{E}_{d\sim\mathcal{D}}-\log\left(\frac{e^{{r_{\theta}(x,y_{\tau(1)})}}}{e^{{r_{\theta}(x,y_{\tau(1)})}}+N_{1}\cdot\mathbb{E}_{y\sim p(y)}[e^{{r_{\theta}(x,y)}}q_{1}(x,y)]}\right),$
$\mathcal{L}_{2}=\mathbb{E}_{d\sim\mathcal{D}}-\log\left(\frac{e^{{r_{\theta}(x,y_{\tau(2)})}}}{e^{{r_{\theta}(x,y_{\tau(2)})}}+N_{2}\cdot\mathbb{E}_{y\sim p(y)}[e^{{r_{\theta}(x,y)}}q_{2}(x,y)]}\right),$
with
$q_{1}(x,y)=\frac{e^{\gamma\cdot r_{est}\left(x,y\right)}}{\sum_{i=2}^{n}e^{\gamma\cdot r_{est}\left(x,y_{\tau(i)}\right)}},$
$q_{2}(x,y)=\frac{e^{\gamma\cdot r_{est}\left(x,y\right)}}{\sum_{i=3}^{n}e^{\gamma\cdot r_{est}\left(x,y_{\tau(i)}\right)}},$
$N_{1}=n-1$, $N_{2}=n-2$, and $p(y)$ is the probability distribution of the dispreferred response $y$. By optimizing the weighted HPS loss $\mathcal{L}_{\boldsymbol{\theta}}$, the model is encouraged to rank $y_{\tau(1)}$ and $y_{\tau(2)}$ above all dispreferred and potentially harmful responses $(y_{\tau(i)})_{i=3}^{n}$, thereby maintaining both helpfulness and response diversity.